OpenStudy (anonymous):

I feel like such a dummy! I cannot figure out how to subtract mixed numbers w borrowing. I am stuck. Any help?

OpenStudy (amistre64):

gotan example to work on?

OpenStudy (amistre64):

and borrowing is fine, the math doesnt mind you borrowing as long as you put it back

OpenStudy (amistre64):

\[-A\frac bc=-A-\frac{b}{c}\]might be useful to know

OpenStudy (phi):

Here's an example \( 4 \frac{1}{2} - 2 \frac{3}{4} \) or, making both fractions have a common denominator \( 4 \frac{2}{4} - 2 \frac{3}{4} \) we can think of this as \( 4+ \frac{2}{4} - (2+ \frac{3}{4}) \) distribute the negative sign, so this is also equal to \( 4+ \frac{2}{4} - 2- \frac{3}{4} \) now if 2/4 were bigger than 3/4 we could just do (4-2) + (2/4-3/4) Actually, we could do that, but that is an ugly answer. Use this idea: 4= 3+1, and 1= 4/4, so 4= 3+ 4/4 and the problem is \( 3+ \frac{4}{4}+\frac{2}{4} - 2- \frac{3}{4} \) switch terms around to get \( 3-2 + \frac{4}{4}+\frac{2}{4} - \frac{3}{4} \) \( (3-2) + (\frac{6}{4} - \frac{3}{4} )\) \( 1 \frac{2}{4} \) or, simplifying the 2/4 \( 1 \frac{1}{2} \)